prove ; sin0(1+tan0)+cos0(1+cot0)=sec0+cosec0

prove ;                                     sin0(1+tan0)+cos0(1+cot0)=sec0+cosec0

Grade:10

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
sin\theta (1+tan\theta ) + cos\theta (1+cot\theta ) = sec\theta +cosec\theta
sin\theta (1+tan\theta ) + cos\theta (1+cot\theta ) = sin\theta (1+tan\theta ) + \frac{cos^{2}\theta}{sin\theta } (1+tan\theta )(1+tan\theta )(\frac{cos^{2}\theta +sin^{2}\theta }{sin\theta }) = \frac{1+tan\theta }{sin\theta } = sec\theta+cosec\theta
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty